OpenStudy (oleg3321):

2+2 super hard please explain will medal

OpenStudy (oleg3321):

please explain how to add 2+2

OpenStudy (anonymous):

2+2 = 4.

OpenStudy (oleg3321):

hahahahah

OpenStudy (xapproachesinfinity):

2+1+1=1+1+2=4 lol